# Speech-FT: Merging Pre-trained And Fine-Tuned Speech Representation Models For Cross-Task Generalization

### v2

@@ Abstract
Index Terms:
speech representation learning, fine-tuning strategy, model merging
@@ I Introduction
Speech representation models have become a cornerstone of modern speech processing [1, 2, 3, 4, 5], offering strong generalization across a wide range of downstream tasks such as phoneme recognition, speaker identification, and emotion recognition [6, 7, 8]. These models learn general speech representations from large-scale unlabeled data via self-supervised pre-training. Recently, there has been growing interest in further enhancing these representations through fine-tuning, either using supervised signals [9] or by continuing the self-supervised pre-training process [10].
Despite its potential benefits, fine-tuning often leads to a degradation in cross-task generalization ability [9]. A fine-tuned model may become overly specialized to the target task, resulting in poor transferability to other tasks. While this issue has been partially attributed to changes in the pre-trained features [11], the underlying causes remain unclear. One common belief is that large deviations in weight space lead to representational drift [12, 13, 14]. However, this is not necessarily the case. Even with substantial weight updates, a model can retain high feature similarity to the original representations, depending on how fine-tuning is conducted [15].
@@
In this work, we investigate how fine-tuning can enhance pre-trained speech representation models while preserving cross-task generalization ability. To achieve this, we propose Speech-FT (Speech Fine-Tuning), a novel two-stage fine-tuning framework designed to maintain cross-task generalization while benefiting from fine-tuning. In the first stage, Speech-FT applies a fine-tuning strategy specifically designed to reduce representational drift with the pre-trained model. We refer to this approach as stable fine-tuning throughout this paper. In the second stage, Speech-FT further restores feature similarity and enhances cross-task generalization through weight-space interpolation between the pre-trained and fine-tuned models. This interpolation step can be interpreted as a form of model merging [16], which has been shown to effectively integrate information from different models. In the context of Speech-FT, interpolation combines task-specific information acquired during fine-tuning with the relatively general representations of the pre-trained model.
As we will analyze in Section V-D, Speech-FT achieves higher feature similarity with the pre-trained model compared to weight-space regularization, despite allowing larger parameter updates during fine-tuning. This aligns with prior work [15] which demonstrates the existence of low-impact subspaces in weight space where substantial weight changes induce minimal functional change; our findings provide empirical evidence suggesting that such invariant subspaces exist in the context of speech representation models as well.
We conduct extensive experiments primarily on HuBERT [4], across a variety of supervised and unsupervised fine-tuning scenarios. In addition, we validate the generality of Speech-FT on other speech representation models including wav2vec 2.0 [2], DeCoAR 2.0 [3], and WavLM Base+ [5]. Our evaluations on the SUPERB benchmark [6] reveal that Speech-FT consistently outperforms other baselines that constrain weight deviation during fine-tuning, such as weight-space regularization and LoRA fine-tuning [17].
@@ II-A Problem Formulation
Let the pre-trained speech representation model be denoted as $\theta_{0}$, whose representations are evaluated on a set of downstream tasks $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$. Given a particular fine-tuning task $\hat{t}$, our objective is to design an algorithm $C$ that produces a fine-tuned model $\hat{\theta}=C(\theta_{0},\hat{t})$ which improves performance across tasks in $\mathcal{T}$.
Figure 1 illustrates the overall pipeline. During fine-tuning, we attach a task prediction model $D$ to the speech representation model $\theta_{0}$. The task prediction model $D$ takes the speech representations generated by $\theta_{0}$ as input and produces task-specific predictions. Both $D$ and $\theta_{0}$ are optimized jointly during fine-tuning. After fine-tuning, the task prediction model $D$ is discarded, and we evaluate the representation of the fine-tuned model $\hat{\theta}$.
Figure 1: Illustration of the pipeline for representation learning and evaluation. Our goal is to find an effective algorithm $C$ to improve the representation of $\hat{\theta}$.
@@
1:  Input: Pre-trained model weights $\theta_{0}$, task prediction model $D$, fine-tuning task $\hat{t}$, scaling factor $\alpha$, stable fine-tuning rate $\beta$, total fine-tuning steps $S$
@@
15:  Let $\theta^{\prime}$ be the model weights after fine-tuning
@@ II-B Speech-FT
Speech-FT consists of two main components: stable fine-tuning and weight-space interpolation. Algorithm 1 provides an overview of the Speech-FT procedure.
@@
When fine-tuning with a randomly initialized task prediction model $D$, the pre-trained features may be severely altered as the model adapts to the fine-tuning task [11]. This issue can be mitigated by first training a well-initialized task prediction model $D$. Therefore, we begin fine-tuning by updating only the task prediction model $D$ during the initial $\beta\%$ of fine-tuning steps.
We further stabilize fine-tuning by freezing the downsampling module. Modern speech representation models often include a downsampling module at the beginning, typically a multi-layer convolutional encoder that processes raw audio and generates downsampled features. Prior studies have shown that this module captures low-level features, such as frequency patterns [18], which are generally crucial for various applications. To preserve these low-level features, we freeze the downsampling module throughout the entire fine-tuning process.
@@
Instead, we adopt model merging as an alternative, combining the pre-trained and fine-tuned models using linear interpolation. Model merging techniques have been shown to effectively integrate information from different models [16, 19, 20, 21, 22]. In our context, interpolating between the pre-trained and fine-tuned models leverages the pre-trained model’s strong cross-task generalization ability while still capturing task-specific improvements. Notably, this approach also helps restore feature similarity with the pre-trained model—an aspect that weight-space regularization often fails to preserve.
@@ II-C Scenarios with Multiple Fine-tuning Tasks
Speech-FT can also be applied to scenarios involving multiple fine-tuning tasks. Given multiple fine-tuning tasks $\{\hat{t_{1}},\hat{t_{2}},\dots,\hat{t_{k}}\}$, our goal is to design an algorithm $C$ such that $\hat{\theta}=C(\theta_{0},\hat{t_{1}},\hat{t_{2}},\dots,\hat{t_{k}})$ leads to improved representations. By leveraging multiple tasks $\hat{t}_{1},\hat{t}_{2},\dots,\hat{t}_{k}$, we aim to incorporate diverse information during fine-tuning, further enhancing speech representations.
@@
Multitask Fine-tuning (MTF). We perform stable fine-tuning on the pre-trained model using multiple tasks simultaneously, where separate task prediction models are appended to the pre-trained model. The resulting multitask model $\theta^{\prime}$ is used in Speech-FT as follows:
@@
TIES Merging. In contrast to linear merging, we adopt a more sophisticated merging technique known as TIES merging [23]. The resulting model in Speech-FT is given by:
@@
TABLE I: A comparison of stable fine-tuning (Stable-FT) and regular fine-tuning.
Fine-tuning Task Method PR SID ER SF $\text{SUPERB}_{\text{S}}\uparrow$
PER%$\downarrow$ ACC%$\uparrow$ ACC%$\uparrow$ F1$\uparrow$ CER%$\downarrow$
SID on Voxceleb1 Stable-FT 29.08 87.75 60.29 77.64 43.27 651.42
Regular fine-tuning 34.68 86.96 56.09 73.73 46.61 568.62
ER on CREMA-D Stable-FT 6.62 81.48 64.69 87.28 27.03 847.33
Regular fine-tuning 6.25 80.1 64.34 86.43 27.97 834.38
@@ III Experiments
All experiments are conducted on HuBERT [4], unless otherwise specified. We evaluate the generality of our method on wav2vec 2.0 [2], DeCoAR 2.0 [3], and WavLM Base+ [5] in Section V-A. Instead of pre-training these models from scratch, we use publicly available pre-trained models [^1] https://github.com/facebookresearch/fairseq [^2] https://github.com/awslabs/speech-representations/ [^3] https://github.com/microsoft/unilm/tree/master/wavlm as the starting point for fine-tuning. HuBERT, wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ contain approximately 94M, 95M, 90M, and 95M parameters, respectively.
For each fine-tuning task, we select the checkpoint with the best validation score. For weight-space interpolation, we fix the scaling factor $\alpha=0.25$ throughout the paper and discuss its impact in Section V-D. For stable fine-tuning, where only the task prediction model $D$ is updated during the initial $\beta\%$ of fine-tuning steps, we set $\beta=10\%$ across all experiments. We fine-tune all model parameters except the downsampling module, which is kept frozen for stable fine-tuning. In Section V-C, we additionally include parameter-efficient fine-tuning baselines for comparison. All fine-tuning experiments are performed on a single NVIDIA RTX 3090 GPU. We outline the hyperparameters, objectives, datasets, task prediction models $D$, and computational budget used for different fine-tuning tasks below.
@@ III-A Supervised Fine-tuning Tasks
Automatic Speech Recognition (ASR) on TED-LIUM. TED-LIUM [24] is a dataset comprising English TED talks paired with corresponding transcriptions. The task prediction model is a three-layer LSTM trained using CTC loss. Fine-tuning is conducted for 50,000 steps with an effective batch size of 32 and a learning rate of $10^{-4}$. We ensure data quality by excluding samples containing the <unk> token and limit the training set to a randomly selected 100-hour subset. The entire process takes approximately 12 hours.
Phoneme Classification (PC) on TIMIT. TIMIT [25] is a widely used English dataset for phoneme classification. The task prediction model is a linear projection layer trained with cross-entropy loss. Fine-tuning is performed for 300,000 steps with a batch size of 16 and a learning rate of $10^{-4}$. Following prior work [26], we use standard training and validation splits, with the training set containing approximately 3.14 hours of data. The entire process takes about 6 hours.
Speaker Identification (SID) on Librispeech.
Librispeech [27] is a corpus of approximately 1,000 hours of read English speech. The task prediction model is a linear projection layer that takes mean-pooled speech representations as input and is trained with cross-entropy loss. Fine-tuning is conducted for 100,000 steps using a batch size of 32 and a learning rate of $2\times 10^{-4}$. Following prior work [28], we randomly sample a 10-hour subset while maintaining the same number of speakers. The entire process takes around 12 hours.
Emotion Recognition (ER) on CREMA-D. CREMA-D [29] is an audio-visual emotional dataset with 7,442 English speech clips from 91 actors. Only speech-based emotional annotations are used. The task prediction model is a linear projection layer that takes mean-pooled speech representations as input and is trained with cross-entropy loss. Fine-tuning is performed for 15,510 steps with a batch size of 32 and a learning rate of $10^{-4}$. Following the EMO-SUPERB setup [30], the training set contains approximately 6.5 hours of speech, and the entire process takes around 5 hours.
@@
Speaker Identification (SID) on VoxCeleb1. VoxCeleb1 [31] is a large-scale speaker identification dataset. The task prediction model is a linear projection layer that processes mean-pooled speech representations and is trained with cross-entropy loss. Fine-tuning is conducted for 200,000 steps with a batch size of 32 and a learning rate of $10^{-4}$, using approximately 318 hours of speech data. The entire process takes around 30 hours.
Emotion Recognition (ER) on IEMOCAP. IEMOCAP [32] is a widely used dataset for emotion recognition. The task prediction model is a linear projection layer that processes mean-pooled speech representations and is trained with cross-entropy loss. Fine-tuning is conducted for 30,000 steps with a batch size of 32 and a learning rate of $10^{-4}$, using Sections 2 to 5 for training and Section 1 for validation. The training set contains approximately 5.56 hours of speech, and the entire process takes around 5 hours.
@@ III-B Unsupervised Fine-tuning Tasks
For unsupervised fine-tuning, we perform continued pre-training of HuBERT on AISHELL-3 [33], a widely used Mandarin speech corpus published by Beijing Shell Shell Technology Co.,Ltd. Our goal is to examine whether Speech-FT enables the model to adapt to a new language (Chinese) while maintaining its performance on a previously learned language (English). We perform continued pre-training for 100,000 steps with an effective batch size of 32 and a learning rate of $10^{-5}$. The entire process takes approximately 24 hours. Since the data in AISHELL-3 has a sampling rate of 44100, we resample it to 16000 to match the sampling rate of HuBERT. The training dataset of AISHELL-3 consist of approximately 63.17 hours of data.
Since the k-means models used to generate the targets for HuBERT Base have never been publicly released (only the model for HuBERT Large is available [^4] https://dl.fbaipublicfiles.com/hubert/hubert_base_ls960_L9_km500.bin), we follow the approach proposed in prior work [34], generating pre-training targets by running HuBERT Base and extracting the $\arg\max$ from its predictions as pseudo targets. We refer to this method as continued pre-training, as it extends the pre-training process of the original model. From another perspective, it can also be regarded as a form of self-training [35], since the pseudo targets are generated using the model itself.
We follow the default behavior of HuBERT [4] when extracting its prediction on targets. Specifically, a feature sequence $[o_{1},\dots,o_{T}]$ is extracted from the encoder. Then, the distribution over codewords is parameterized with
$\mathbf{P}_{t,c}=\frac{\exp\left(\operatorname{sim}\left(Ao_{t},e_{c}\right)/\tau\right)}{\sum_{c^{\prime}=1}^{C}\exp\left(\operatorname{sim}\left(Ao_{t},e_{c^{\prime}}\right)/\tau\right)},$ (5)
where $A$ is the projection matrix, $e_{c}$ is the embedding for codeword $c$, $\operatorname{sim}(\cdot,\cdot)$ computes the cosine similarity between two vectors, and $\tau$ scales the logit, which is set to 0.1. Then, the pseudo targets $\mathbf{y}=[y_{1},\dots,y_{T}]$ are then obtained by applying the $\arg\max$:
$y_{t}=\arg\max_{c}\mathbf{P}_{t,c},\quad\forall t\in\{1,\dots,T\}$ (6)
Although the dataset used for continued pre-training is a Mandarin speech corpus, we found that this method still generates a reasonable sequence of pseudo targets without significant cluster imbalance issues.
We adopt a simplified variant of the HuBERT loss as our continued pre-training objective [18], which is defined as a cross-entropy loss between the pseudo targets $y_{t}$ and the output of a linear classifier at masked time steps. Specifically, we train a new projection matrix $W$ using the following objective:
$\mathcal{L}=\frac{1}{|M|}\sum_{t\in M}\text{CE}(Wo_{t},y_{t})$ (7)
Here, $M$ represents the set of masked time steps. The function $\text{CE}(\cdot,\cdot)$ refers to the cross-entropy loss. We follow the masking configuration of the original HuBERT. This variant of the HuBERT loss has been shown to be as effective as the original version [18], while also being more computationally efficient [36]. Furthermore, optimization without a learnable codebook is relatively stable, making it a suitable choice for continued pre-training.
TABLE II: SUPERB score results on eight supervised fine-tuning tasks.
Method ASR PC SID ER ASR PR SID ER
(TED-LIUM) (TIMIT) (Librispeech) (CREMA-D) (Librispeech) (Librispeech) (Voxceleb1) (IEMOCAP)
Speech-FT 905.79 877.66 877.97 871.93 882.44 883.93 881.07 1010.29
Stable-FT 870.71 726.64 847.33 741.19 805.33 842.03 651.42 908.86
Pre-trained 870.20
TABLE III: Detailed SUPERB downstream results for eight supervised fine-tuning tasks. Background color highlights cases where the fine-tuning task matches the evaluation task.
Fine-tuning Task Method PR on SID on ER on SF on
@@
PER%$\downarrow$ ACC%$\uparrow$ ACC%$\uparrow$ F1$\uparrow$ CER%$\downarrow$
ASR on TED-LIUM Speech-FT 3.94 84.11 67.78 88.84 23.46
Stable-FT 5.22 79.89 65.38 88.52 23.80
PC on TIMIT Speech-FT 4.76 81.78 65.48 88.65 24.05
Stable-FT 10.34 66.34 61.25 83.50 33.82
SID on Librispeech Speech-FT 5.67 83.24 65.48 87.98 24.90
Stable-FT 14.18 75.52 59.51 84.73 31.54
ER on CREMA-D Speech-FT 5.02 83.36 65.59 88.41 24.75
Stable-FT 6.62 81.48 64.69 87.28 27.03
ASR on Librispeech Speech-FT 3.93 82.79 66.38 87.97 24.93
Stable-FT 5.35 72.14 63.82 86.01 29.32
PR on Librispeech Speech-FT 2.99 81.95 65.79 88.60 24.52
Stable-FT 2.21 75.60 63.14 88.09 26.24
SID on Voxceleb1 Speech-FT 6.91 88.80 66.03 87.20 25.23
Stable-FT 29.08 87.75 60.29 77.64 43.27
ER on IEMOCAP Speech-FT 4.95 82.63 85.95 87.23 25.80
Stable-FT 16.11 73.07 90.07 81.40 36.62
Pre-trained 5.17 81.86 64.99 88.54 24.70

### v3

@@ I Introduction
Speech representation models have become a cornerstone of modern speech processing [baevski2020wav2vec, ling2020decoar, hsu2021hubert, chen2022wavlm], offering strong generalization across a wide range of downstream tasks such as phoneme recognition, speaker identification, and emotion recognition [yang2021superb]. These models learn general speech representations from large-scale unlabeled data via self-supervised pre-training. Recently, there has been growing interest in further enhancing these representations through fine-tuning, either using supervised signals [chen2021speech] or by continuing the self-supervised pre-training process [getman2024happens].
Despite its potential benefits, fine-tuning often leads to a degradation in cross-task generalization ability [chen2021speech]. A fine-tuned model may become overly specialized to the target task, resulting in poor transferability to other tasks. While this issue has been partially attributed to changes in the pre-trained features [kumar2022fine], the underlying causes remain unclear. One common belief is that large deviations in weight space lead to representational drift [kirkpatrick2017overcoming, xu2019forget, chen-etal-2020-recall]. However, this is not necessarily the case. Even with substantial weight updates, a model can retain high feature similarity to the original representations, depending on how fine-tuning is conducted [raghavan2024engineering].
@@
In this work, we investigate how fine-tuning can enhance pre-trained speech representation models while preserving cross-task generalization ability. To achieve this, we propose Speech-FT (Speech Fine-Tuning), a novel two-stage fine-tuning framework designed to maintain cross-task generalization while benefiting from fine-tuning. In the first stage, Speech-FT applies a fine-tuning strategy specifically designed to reduce representational drift with the pre-trained model. We refer to this approach as stable fine-tuning throughout this paper. In the second stage, Speech-FT further restores feature similarity and enhances cross-task generalization through weight-space interpolation between the pre-trained and fine-tuned models. This interpolation step can be interpreted as a form of model merging [ilharco2022editing], which has been shown to effectively integrate information from different models. In the context of Speech-FT, interpolation combines task-specific information acquired during fine-tuning with the relatively general representations of the pre-trained model.
As we will analyze in Section V-D, Speech-FT achieves higher feature similarity with the pre-trained model compared to weight-space regularization, despite allowing larger parameter updates during fine-tuning. This aligns with prior work [raghavan2024engineering] which demonstrates the existence of low-impact subspaces in weight space where substantial weight changes induce minimal functional change; our findings provide empirical evidence suggesting that such invariant subspaces exist in the context of speech representation models as well.
We conduct extensive experiments primarily on HuBERT [hsu2021hubert], across a variety of supervised and unsupervised fine-tuning scenarios. In addition, we validate the generality of Speech-FT on other speech representation models including wav2vec 2.0 [baevski2020wav2vec], DeCoAR 2.0 [ling2020decoar], and WavLM Base+ [chen2022wavlm]. Our evaluations on the SUPERB benchmark [yang2021superb] reveal that Speech-FT consistently outperforms other baselines that constrain weight deviation during fine-tuning, such as weight-space regularization and LoRA fine-tuning [hu2022lora].
@@ II-A Problem Formulation
Let the pre-trained speech representation model be denoted as $\theta_{0}$, whose representations are evaluated on a set of downstream tasks $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$. Given a particular fine-tuning task $\hat{t}$, our objective is to design a fine-tuning algorithm that produces a representation model $\hat{\theta}$ which improves performance across tasks in $\mathcal{T}$.
During fine-tuning, we attach a task prediction model $D$ to the speech representation model. The task prediction model $D$ takes the speech representations as input and produces task-specific predictions. Both $D$ and $\theta_{0}$ are optimized jointly during fine-tuning. After fine-tuning, the tuned task prediction model $D$ is discarded (Step 1 in Figure 1). It is worth noting that we evaluate the cross-task performance of the representation model, rather than that of each tuned task prediction model $D$. To examine whether the representation model preserves its cross-task generalization ability, we adopt the SUPERB framework [yang2021superb] (Step 3 in Figure 1), where downstream models are re-tuned individually without relying on the task prediction models trained during the initial fine-tuning stage.
Figure 1: The pipeline of Speech-FT for representation learning and evaluation. Step 1: A pre-trained representation model $\theta_{0}$ undergoes stable fine-tuning on a specific task $\hat{t}$, producing a tuned representation model $\theta^{\prime}$ while discarding the task prediction model $D$. Step 2: The pre-trained and tuned models are merged in weight space to obtain the final representation model $\hat{\theta}$. Step 3: The merged model is evaluated on the SUPERB benchmark by re-training task-specific downstream models, ensuring that cross-task generalization of $\hat{\theta}$ is measured rather than performance of the discarded $D$.
@@
1:Input: Pre-trained model weights $\theta_{0}$, task prediction model $D$, fine-tuning task $\hat{t}$, interpolation scaling factor $\alpha$, stable fine-tuning rate $\beta$, total fine-tuning steps $S$
@@
15: Let $\theta^{\prime}$ be the model weights after stable fine-tuning
@@ II-B Speech-FT
Speech-FT consists of two main components: stable fine-tuning and weight-space interpolation. Algorithm 1 and Steps 1 and 2 of Figure 1 provide an overview of the Speech-FT procedure.
@@
When fine-tuning with a randomly initialized task prediction model $D$, the pre-trained features may be severely altered as the model adapts to the fine-tuning task [kumar2022fine]. This issue can be mitigated by first training a well-initialized task prediction model $D$. Therefore, we begin fine-tuning by updating only the task prediction model $D$ during the initial $\beta\%$ of fine-tuning steps.
We further stabilize fine-tuning by freezing the downsampling module. Modern speech representation models often include a downsampling module at the beginning, typically a multi-layer convolutional encoder that processes raw audio and generates downsampled features. Prior studies have shown that this module captures low-level features, such as frequency patterns [lin2023melhubert], which are generally crucial for various applications. To preserve these low-level features, we freeze the downsampling module throughout the entire fine-tuning process.
@@
Instead, we adopt model merging as an alternative, combining the pre-trained and fine-tuned models using linear interpolation. Model merging techniques have been shown to effectively integrate information from different models [ilharco2022editing, wortsman2022robust, lin2023spurious, lin2024mitigating, kulshreshtha2024sequential]. In our context, interpolating between the pre-trained and fine-tuned models leverages the pre-trained model’s strong cross-task generalization ability while still capturing task-specific improvements. Notably, this approach also helps restore feature similarity with the pre-trained model—an aspect that weight-space regularization often fails to preserve.
@@ II-C Scenarios with Multiple Fine-tuning Tasks
Speech-FT can also be applied to scenarios involving multiple fine-tuning tasks. By leveraging multiple fine-tuning tasks $\hat{t}_{1},\hat{t}_{2},\dots,\hat{t}_{k}$, we aim to incorporate diverse information during fine-tuning, further enhancing speech representations.
@@
Multitask Fine-tuning (MTF). We perform stable fine-tuning on the pre-trained model using multiple tasks simultaneously, where separate task prediction models are appended to the pre-trained model. The resulting multitask model $\theta^{\prime}$ is employed in Speech-FT as described in Equation 1.
@@
TIES Merging. In contrast to linear merging, we adopt a more sophisticated merging technique known as TIES merging [yadav2024ties]. TIES-Merging addresses parameter interference across fine-tuned models by performing three key steps: (i) trimming redundant parameter updates (those minimally changed during fine-tuning), (ii) resolving sign conflicts in parameter directions, and (iii) merging only the parameters whose sign agrees with the chosen consensus sign. The resulting model in Speech-FT is given by:
@@ III Experiments
All experiments are conducted on HuBERT [hsu2021hubert], unless otherwise specified. We evaluate the generality of our method on wav2vec 2.0 [baevski2020wav2vec], DeCoAR 2.0 [ling2020decoar], and WavLM Base+ [chen2022wavlm] in Section V-A. HuBERT, wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ contain approximately 94M, 95M, 90M, and 95M parameters, respectively. Instead of pre-training these models from scratch, we use publicly available pre-trained models [^1] https://github.com/facebookresearch/fairseq [^2] https://github.com/awslabs/speech-representations/ [^3] https://github.com/microsoft/unilm/tree/master/wavlm as the starting point for fine-tuning. While fine-tuning inherently requires additional task-specific labeled data beyond the pre-training stage, all fine-tuning methods in this work are trained on exactly the same datasets and splits, ensuring fully fair comparisons.
For each fine-tuning task, we select the checkpoint with the best validation score. For weight-space interpolation, we fix the scaling factor $\alpha=0.25$ throughout the paper and discuss its impact in Section V-D. We observe that $\alpha=0.25$ generalizes well across different fine-tuning tasks. For stable fine-tuning, where only the task prediction model $D$ is updated during the initial $\beta\%$ of fine-tuning steps, we set $\beta=10\%$ across all experiments, as preliminary experiments showed that a short warm-up is sufficient. We fine-tune all model parameters except the downsampling module, which is kept frozen for stable fine-tuning. In Section V-C, we additionally include parameter-efficient fine-tuning baselines for comparison. All fine-tuning experiments are performed on a single NVIDIA RTX 3090 GPU. We outline the hyperparameters, objectives, datasets, task prediction models $D$, and computational budget used for different fine-tuning tasks below.
@@ III-A Supervised Fine-tuning Tasks
Automatic Speech Recognition (ASR) on TED-LIUM [rousseau2012ted]. The task prediction model is a three-layer LSTM trained using CTC loss. Fine-tuning is conducted for 50,000 steps with an effective batch size of 32 and a learning rate of $10^{-4}$. We limit the training set to a randomly selected 100-hour subset. The entire process takes approximately 12 hours.
Phoneme Classification (PC) on TIMIT [garofolo1993darpa]. The task prediction model is a linear projection layer trained with cross-entropy loss. Fine-tuning is performed for 300,000 steps with a batch size of 16 and a learning rate of $10^{-4}$. Following prior work [lopes2011phone], we use standard training and validation splits. The entire process takes about 6 hours.
Speaker Identification (SID) on Librispeech [panayotov2015librispeech].
The task prediction model is a linear projection layer that takes mean-pooled speech representations as input and is trained with cross-entropy loss. Fine-tuning is conducted for 100,000 steps using a batch size of 32 and a learning rate of $2\times 10^{-4}$. Following prior work [oord2018representation], we randomly sample a 10-hour subset while maintaining the same number of speakers. The entire process takes around 12 hours.
Emotion Recognition (ER) on CREMA-D [cao2014crema]. CREMA-D is an audio-visual emotional dataset. Only speech-based emotional annotations are used. The task prediction model is a linear projection layer that takes mean-pooled speech representations as input and is trained with cross-entropy loss. Fine-tuning is performed for 15,510 steps with a batch size of 32 and a learning rate of $10^{-4}$. Following the EMO-SUPERB setup [wu2024emo], the training set contains approximately 6.5 hours of speech, and the entire process takes around 5 hours.
@@
Speaker Identification (SID) on VoxCeleb1 [nagrani2020voxceleb]. The task prediction model is a linear projection layer that processes mean-pooled speech representations and is trained with cross-entropy loss. Fine-tuning is conducted for 200,000 steps with a batch size of 32 and a learning rate of $10^{-4}$, using approximately 318 hours of speech data. The entire process takes around 30 hours.
Emotion Recognition (ER) on IEMOCAP [busso2008iemocap]. The task prediction model is a linear projection layer that processes mean-pooled speech representations and is trained with cross-entropy loss. Fine-tuning is conducted for 30,000 steps with a batch size of 32 and a learning rate of $10^{-4}$, using Sections 2 to 5 for training and Section 1 for validation. The training set contains approximately 5.56 hours of speech, and the entire process takes around 5 hours.
@@ III-B Unsupervised Fine-tuning Tasks
For unsupervised fine-tuning, we perform continued pre-training of HuBERT on AISHELL-3 [shi2020aishell], a widely used Mandarin speech corpus published by Beijing Shell Shell Technology Co.,Ltd. Our goal is to examine whether Speech-FT enables the model to adapt to a new language (Chinese) while maintaining its performance on a previously learned language (English). We perform continued pre-training for 100,000 steps with an effective batch size of 32 and a learning rate of $10^{-5}$. The entire process takes approximately 24 hours. Since the data in AISHELL-3 has a sampling rate of 44100, we resample it to 16000 to match the sampling rate of HuBERT. The training dataset of AISHELL-3 consist of approximately 63.17 hours of data.
Since the k-means model used to generate the targets for HuBERT Base is not publicly available, we follow prior work [lin2024daisy] and use HuBERT Base’s own predictions (via $\arg\max$ over codeword logits) as pseudo targets for continued pre-training. We then optimize a simplified HuBERT loss [lin2023melhubert, yang2023fast] using cross-entropy over masked time steps, which avoids the need for a learnable codebook and offers stable and efficient training. We refer to this method as continued pre-training, as it extends the pre-training process of the original model. From another perspective, it can also be regarded as a form of self-training [kahn2020self], since the pseudo targets are generated using the model itself.
TABLE I: SUPERB downstream results for eight supervised fine-tuning tasks. Background color highlights cases where the fine-tuning task matches the evaluation task.
Fine-tuning Task Method PR on SID on ER on SF on $\text{SUPERB}_{\text{S}}\uparrow$
@@
[…]
ASR on TED-LIUM Speech-FT 3.94 84.11 67.78 88.84 23.46 905.79
Stable-FT 5.22 79.89 65.38 88.52 23.80 870.71
PC on TIMIT Speech-FT 4.76 81.78 65.48 88.65 24.05 877.66
Stable-FT 10.34 66.34 61.25 83.50 33.82 726.64
SID on Librispeech Speech-FT 5.67 83.24 65.48 87.98 24.90 877.97
Stable-FT 14.18 75.52 59.51 84.73 31.54 847.33
ER on CREMA-D Speech-FT 5.02 83.36 65.59 88.41 24.75 871.93
Stable-FT 6.62 81.48 64.69 87.28 27.03 741.19
ASR on Librispeech Speech-FT 3.93 82.79 66.38 87.97 24.93 882.44
Stable-FT 5.35 72.14 63.82 86.01 29.32 805.33
PR on Librispeech Speech-FT 2.99 81.95 65.79 88.60 24.52 883.93
Stable-FT 2.21 75.60 63.14 88.09 26.24 842.03
SID on Voxceleb1 Speech-FT 6.91 88.80 66.03 87.20 25.23 881.07
Stable-FT 29.08 87.75 60.29 77.64 43.27 651.42
ER on IEMOCAP Speech-FT 4.95 82.63 85.95 87.23 25.80 1010.29
Stable-FT 16.11 73.07 90.07 81.40 36.62 908.86
Pre-trained 5.17 81.86 64.99 88.54 24.70 870.20